40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 186     Entry time: Tue Jun 29 17:57:40 2010
Author: tarac 
Type: Notes 
Category: PMC 
Subject: coupling eff vs Temperature 

We had a problem with power coupling efficiency to the PMC.  Sometime it changes from 80% to 40% or less.

When the temperature of the NPRO is varied by a voltage calibrator @ slow actuator input of the laser controller,

at certain T(V), the coupling efficiency drops from 80% to 40% or less. This might be mode hopping.  However, the transmitted beam power still oscillates ,

and one can see the beam spot pulsating on the screen. I thought it might be mode hopping too, but after changing T, the fluctuation of the power has not gone, there is some suppression though. So it might be the servo. I'll check the TF of the servo and compare it with the schematic, D980352-B-1.

ELOG V3.1.3-